selling price per unit = 450
variable expense per unit = 280
total fixed expense = 2,000,000

1. calculate the accounting breakeven point in $?

2.what is the degree of operating leverage if current sales are 16,000 units

how would you interpret your findings
Answer #1

Part 1

Breakeven point in dollars = fixed expenses / contribution margin ratio

Contribution margin ratio = (selling price per unit – variable cost per unit)/selling price per unit = (450-280)/450 = 37.78%

Breakeven point in dollars = 2000000/37.78%= $5293806 (rounding contribution margin ratio to two decimal places)

If not it is not to be rounded then it will be

Breakeven point in dollars = 2000000/((450-280)/450) = $5294118

Part 2

Sales (450*16000)

7200000

Less: variable expense (280*16000)

4480000

Contribution margin

2720000

Less: fixed expenses

2000000

Net operating income

720000

Degree of operating leverage = contribution margin / net operating income = 2720000/720000 = 3.78

Part 3

The degree of operating leverage measures the change in the operating income with respect to a change in sales. If there is a large proportion of fixed costs compared to variable costs then there will be higher levels of operating leverage. Here, the proportion of variable cost is higher than fixed cost and therefore the level of operating leverage can be considered low. This means it has a smaller profit on each sale, and is not required to increase sales to cover its lower fixed expenses.